Two angles of an isosceles triangle are x and (x+10).find two possible value of x.
Answer by TimothyLamb(4379)

case 1:
x + x + (x+10) = 180
3x + 10 = 180
3x = 170
x = 56.666666 degrees
---
case 2:
x + (x+10) + (x+10) = 180
3x + 20 = 180
3x = 160
x = 53.333333 degrees
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php